Haemolytic jaundice is due to dominant gene but only 20% of the people develop this disease. A heterozygous man marries a homozygous normal woman. What proportion of the children in population would be expected to have this disorder and how?
(1) 1/5
(2) 1/20
(3) 1/10
(4) 1/2
1
Expert's answer
2015-09-08T07:00:31-0400
Man: Aa. Women: aa.
Children: Aa, Aa, aa, aa. Aa have disorder, aa – does not.
The proportion of the children in population would be expected to have this disorder: 2/4*20%=1/2*1/5=1/10.